OpenStudy (meilendurcer):

1. Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 2. Use the graph below to fill in the blank with the correct number. f(1) = _______ 3. Let f(p) be the average number of days a house stays on the market before being sold for price p in $1,000s. Which statement best describes the meaning of f(250)? A. The house sold for $250,000. B. The house stayed on the market for an average of 250 days before being sold. C. This is the average number of days the house stayed on the market before being sold for $250,000.

OpenStudy (meilendurcer):

D. The house sold on the market for $250,000 and stayed on the market for an average of 250 days before being sold.

OpenStudy (anonymous):

1) 0,1,1,2,3,5,8... 3) Option B

OpenStudy (meilendurcer):

i've done this test before and number 3 isn't the right answer

OpenStudy (meilendurcer):

i mean the answer for number 3

OpenStudy (meilendurcer):

i was thinking D but i'm not too sure on that one

OpenStudy (anonymous):

Its option C)This is the average number of days the house stayed on the market before being sold for $250,000 ....I mistakenly choose B..I am sorry

OpenStudy (meilendurcer):

it's fine i did too. is the answer for number 2, two?

OpenStudy (anonymous):

I didn't understand your 2nd question.. I think so its incomplete..

OpenStudy (meilendurcer):

oh the second question is to find f(1) from the graph i added

OpenStudy (meilendurcer):

f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

Maybe answer is -2

OpenStudy (meilendurcer):

for the last one or number 2?

OpenStudy (anonymous):

f(1)=-2

OpenStudy (meilendurcer):

are you sure it isn't positive 2?

OpenStudy (anonymous):

-2 is the only answer I can make out of...

OpenStudy (meilendurcer):

ok :)

OpenStudy (meilendurcer):

do you happen to know the other one? from the formulas i think it's 10

OpenStudy (anonymous):

Which one??

OpenStudy (meilendurcer):

f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

22...I think so..

OpenStudy (meilendurcer):

oh wait yeah i think it is. I remember it now. yay 9 out of 10 thanks